# Flock: A Knowledge Graph Foundation Model via Learning on Random Walks

### v1

@@ Results.
We report the average zero-shot performance for entity prediction and relation prediction over 54 KGs in Tables 7(a) and 7(b), respectively. Across all tasks, Trix with naive noise injection fails to close the gap between Flock. In particular, Trix + noise degrades compared with vanilla Trix without noise injection in relation prediction, while boosting the performance in the entity prediction task. We hypothesize that such a difference lies in the added randomization breaks symmetry among entity embeddings more than among relation embeddings, and entity prediction depends more on having distinguishable entity representations than relation prediction does. Additionally, we attribute this performance gap between Flock and Trix + noise to the source of randomization. Flock introduces stochasticity through random walks, which induces structure-informed perturbations that respect the underlying topology. In contrast, Trix with naive noise injection attempts to break deterministic node-relation equivariance by introducing structure-agonistic noise, which might, in turn, hurt the model’s generalization. Together, these findings suggest that simply adding structure-agonistic noise is insufficient; performance gains only arise when stochasticity is topology-aware and is induced from the graph structure in a principled way.
Appendix G Further experimental details
Table 8: Zero-shot inductive link prediction results (MRR and Hits@10) for Ultra, Trix, and Flock. Bold indicates the best score per row.
Dataset Ultra Trix Flock
MRR Hits@10 MRR Hits@10 MRR Hits@10
Inductive $e,r$
FB-25 0.388 0.640 0.393 0.650 0.404 0.664
FB-50 0.338 0.543 0.334 0.547 0.352 0.566
FB-75 0.403 0.604 0.401 0.611 0.418 0.622
FB-100 0.449 0.642 0.436 0.635 0.452 0.663
WK-25 0.316 0.532 0.305 0.496 0.280 0.491
WK-50 0.166 0.324 0.166 0.313 0.136 0.278
WK-75 0.365 0.537 0.368 0.513 0.382 0.538
WK-100 0.164 0.286 0.188 0.299 0.187 0.304
NL-0 0.342 0.523 0.385 0.549 0.381 0.606
NL-25 0.395 0.569 0.377 0.589 0.345 0.590
NL-50 0.407 0.570 0.404 0.548 0.366 0.565
NL-75 0.368 0.547 0.351 0.525 0.311 0.524
NL-100 0.471 0.651 0.486 0.676 0.452 0.692
MT1 tax 0.224 0.305 0.358 0.452 0.282 0.383
MT1 health 0.298 0.374 0.376 0.457 0.385 0.481
MT2 org 0.095 0.159 0.091 0.156 0.100 0.163
MT2 sci 0.258 0.354 0.323 0.465 0.318 0.458
MT3 art 0.259 0.402 0.284 0.441 0.301 0.466
MT3 infra 0.619 0.755 0.655 0.797 0.684 0.821
MT4 sci 0.274 0.449 0.290 0.460 0.301 0.463
MT4 health 0.624 0.737 0.677 0.775 0.680 0.780
Metafam 0.238 0.644 0.341 0.815 0.476 0.935
FBNELL 0.485 0.652 0.473 0.660 0.502 0.700
Inductive $e$
FB-v1 0.498 0.656 0.515 0.682 0.500 0.697
FB-v2 0.512 0.700 0.525 0.730 0.535 0.737
FB-v3 0.491 0.654 0.501 0.669 0.511 0.685
FB-v4 0.486 0.677 0.493 0.687 0.505 0.702
WN-v1 0.648 0.768 0.699 0.791 0.698 0.803
WN-v2 0.663 0.765 0.678 0.781 0.696 0.790
WN-v3 0.376 0.476 0.418 0.541 0.467 0.608
WN-v4 0.611 0.705 0.648 0.723 0.653 0.729
NL-v1 0.785 0.913 0.806 0.898 0.658 0.863
NL-v2 0.526 0.707 0.569 0.768 0.588 0.797
NL-v3 0.515 0.702 0.558 0.743 0.590 0.783
NL-v4 0.479 0.712 0.538 0.765 0.555 0.786
HM 1k 0.059 0.092 0.072 0.128 0.069 0.119
HM 3k 0.037 0.077 0.069 0.119 0.067 0.118
HM 5k 0.034 0.071 0.062 0.110 0.064 0.116
HM Indigo 0.440 0.648 0.436 0.645 0.423 0.638
ILPC Small 0.302 0.443 0.303 0.455 0.309 0.459
ILPC Large 0.290 0.424 0.307 0.428 0.318 0.438
Table 9: Zero-shot transductive link prediction results (MRR and Hits@10) for Ultra, Trix, and Flock. Bold indicates the best score per row.
Dataset Ultra Trix Flock
MRR Hits@10 MRR Hits@10 MRR Hits@10
Transductive
NELL995 0.406 0.543 0.472 0.629 0.494 0.655
NELL23k 0.239 0.408 0.290 0.497 0.233 0.398
WDsinger 0.382 0.498 0.511 0.609 0.410 0.528
ConceptNet100k 0.082 0.162 0.193 0.345 0.248 0.453
CoDEx Small 0.472 0.667 0.472 0.670 0.441 0.644
CoDEx Large 0.338 0.469 0.335 0.469 0.342 0.464
YAGO310 0.451 0.615 0.409 0.627 0.414 0.674
AristoV4 0.182 0.282 0.181 0.286 0.308 0.443
DBpedia100k 0.398 0.576 0.426 0.603 0.450 0.627
Hetionet 0.257 0.379 0.279 0.420 0.246 0.371
FB15k-237(10) 0.248 0.398 0.246 0.393 0.246 0.402
FB15k-237(20) 0.272 0.436 0.269 0.430 0.273 0.444
FB15k-237(50) 0.324 0.526 0.321 0.521 0.319 0.518
Table 10: Finetuned inductive link prediction results (MRR and Hits@10) for Ultra, Trix, and Flock. Bold indicates the best score per row.
Dataset Ultra Trix Flock
MRR Hits@10 MRR Hits@10 MRR Hits@10
Inductive $e,r$
FB-25 0.383 0.635 0.393 0.650 0.405 0.666
FB-50 0.334 0.538 0.334 0.547 0.357 0.570
FB-75 0.400 0.598 0.401 0.611 0.425 0.630
FB-100 0.444 0.643 0.436 0.633 0.460 0.668
WK-25 0.321 0.535 0.300 0.493 0.298 0.506
WK-50 0.140 0.280 0.166 0.313 0.127 0.260
WK-75 0.380 0.530 0.368 0.513 0.405 0.556
WK-100 0.168 0.286 0.188 0.299 0.187 0.306
NL-0 0.329 0.551 0.385 0.549 0.418 0.619
NL-25 0.407 0.596 0.377 0.589 0.405 0.626
NL-50 0.418 0.595 0.405 0.555 0.391 0.562
NL-75 0.374 0.570 0.351 0.525 0.344 0.544
NL-100 0.458 0.684 0.482 0.691 0.486 0.714
MT1 tax 0.330 0.459 0.397 0.508 0.413 0.497
MT1 health 0.380 0.467 0.376 0.457 0.394 0.493
MT2 org 0.104 0.170 0.098 0.162 0.107 0.174
MT2 sci 0.311 0.451 0.331 0.526 0.366 0.525
MT3 art 0.306 0.473 0.289 0.441 0.330 0.483
MT3 infra 0.657 0.807 0.672 0.810 0.709 0.838
MT4 sci 0.303 0.478 0.305 0.482 0.324 0.509
MT4 health 0.704 0.785 0.702 0.785 0.711 0.790
Metafam 0.997 1.000 0.997 1.000 0.992 1.000
FBNELL 0.481 0.661 0.478 0.655 0.531 0.714
Inductive $e$
FB-v1 0.509 0.670 0.515 0.682 0.549 0.721
FB-v2 0.524 0.710 0.525 0.730 0.553 0.754
FB-v3 0.504 0.663 0.501 0.669 0.528 0.696
FB-v4 0.496 0.684 0.493 0.687 0.510 0.702
WN-v1 0.685 0.793 0.705 0.798 0.715 0.811
WN-v2 0.679 0.779 0.682 0.780 0.702 0.795
WN-v3 0.411 0.546 0.425 0.543 0.494 0.627
WN-v4 0.614 0.720 0.650 0.722 0.665 0.741
NL-v1 0.757 0.878 0.804 0.899 0.762 0.928
NL-v2 0.575 0.761 0.571 0.764 0.612 0.806
NL-v3 0.563 0.755 0.571 0.759 0.606 0.803
NL-v4 0.469 0.733 0.551 0.772 0.572 0.801
HM 1k 0.042 0.100 0.072 0.128 0.071 0.153
HM 3k 0.030 0.090 0.069 0.119 0.067 0.153
HM 5k 0.025 0.068 0.074 0.118 0.061 0.130
HM Indigo 0.432 0.639 0.436 0.645 0.418 0.633
ILPC Small 0.303 0.453 0.303 0.455 0.305 0.454
ILPC Large 0.308 0.431 0.310 0.431 0.320 0.441
Table 11: Finetune transductive link prediction results (MRR and Hits@10) for Ultra, Trix, and Flock. Bold indicates the best score per row.
Dataset Ultra Trix Flock
MRR Hits@10 MRR Hits@10 MRR Hits@10
Pretrained
FB15k-237 0.368 0.564 0.366 0.559 0.343 0.532
WN18RR 0.480 0.614 0.514 0.611 0.550 0.656
CoDEx Medium 0.372 0.525 0.365 0.521 0.351 0.496
Transductive
NELL995 0.509 0.660 0.506 0.648 0.531 0.665
NELL23k 0.268 0.450 0.306 0.536 0.280 0.465
WDsinger 0.417 0.526 0.502 0.620 0.435 0.543
ConceptNet100k 0.310 0.529 0.340 0.564 0.352 0.580
CoDEx Small 0.490 0.686 0.484 0.676 0.463 0.648
CoDEx Large 0.343 0.478 0.348 0.481 0.342 0.467
YAGO310 0.557 0.710 0.541 0.702 0.552 0.700
AristoV4 0.343 0.496 0.345 0.499 0.383 0.523
DBpedia100k 0.436 0.603 0.457 0.619 0.470 0.623
Hetionet 0.399 0.538 0.394 0.534 0.314 0.465
FB15k-237(10) 0.254 0.411 0.253 0.408 0.260 0.420
FB15k-237(20) 0.274 0.445 0.273 0.441 0.284 0.459
FB15k-237(50) 0.325 0.528 0.322 0.522 0.317 0.517
Table 12: Zero-shot inductive relation prediction results (MRR and Hits@1) for Ultra, Trix, and Flock. Bold indicates the best score per row.
Dataset Ultra Trix Flock
MRR Hits@1 MRR Hits@1 MRR Hits@1
Inductive $e,r$
FB-25 0.687 0.565 0.805 0.724 0.895 0.839
FB-50 0.696 0.575 0.780 0.699 0.880 0.820
FB-75 0.698 0.555 0.822 0.747 0.903 0.844
FB-100 0.830 0.728 0.921 0.880 0.962 0.938
WK-25 0.857 0.760 0.881 0.823 0.952 0.929
WK-50 0.865 0.793 0.868 0.818 0.921 0.882
WK-75 0.911 0.875 0.916 0.883 0.962 0.944
WK-100 0.887 0.812 0.907 0.869 0.963 0.937
NL-0 0.632 0.502 0.658 0.519 0.714 0.574
NL-25 0.688 0.562 0.742 0.614 0.729 0.632
NL-50 0.680 0.569 0.755 0.636 0.813 0.728
NL-75 0.795 0.692 0.788 0.699 0.833 0.756
NL-100 0.743 0.564 0.884 0.796 0.939 0.889
MT1 tax 0.985 0.976 0.975 0.958 0.998 0.997
MT1 health 0.721 0.561 0.973 0.949 0.991 0.983
MT2 org 0.974 0.951 0.986 0.973 0.991 0.984
MT2 sci 0.976 0.961 0.964 0.941 0.995 0.992
MT3 art 0.881 0.798 0.885 0.825 0.944 0.907
MT3 infra 0.962 0.935 0.940 0.905 0.989 0.980
MT4 sci 0.933 0.891 0.966 0.944 0.974 0.957
MT4 health 0.826 0.719 0.937 0.898 0.990 0.983
Metafam 0.124 0.000 0.291 0.011 0.490 0.223
FBNELL 0.700 0.564 0.726 0.605 0.833 0.737
Inductive $e$
FB-v1 0.646 0.523 0.705 0.599 0.814 0.723
FB-v2 0.695 0.570 0.713 0.590 0.847 0.761
FB-v3 0.679 0.553 0.742 0.644 0.860 0.780
FB-v4 0.638 0.488 0.766 0.665 0.873 0.799
WN-v1 0.836 0.740 0.792 0.613 0.924 0.858
WN-v2 0.853 0.790 0.764 0.572 0.924 0.863
WN-v3 0.707 0.577 0.741 0.568 0.937 0.888
WN-v4 0.860 0.803 0.764 0.570 0.937 0.886
NL-v1 0.636 0.358 0.657 0.453 0.862 0.731
NL-v2 0.742 0.652 0.780 0.696 0.893 0.855
NL-v3 0.669 0.544 0.725 0.612 0.815 0.731
NL-v4 0.606 0.489 0.794 0.691 0.868 0.807
ILPC Small 0.905 0.843 0.919 0.872 0.955 0.921
ILPC Large 0.875 0.799 0.894 0.829 0.948 0.908
HM 1k 0.626 0.447 0.663 0.414 0.687 0.500
HM 3k 0.592 0.439 0.664 0.418 0.714 0.549
HM 5k 0.605 0.452 0.672 0.428 0.746 0.593
HM Indigo 0.681 0.559 0.852 0.765 0.956 0.921
Table 13: Zero-shot transductive relation prediction results (MRR and Hits@1) for Ultra, Trix, and Flock.
Dataset Ultra Trix Flock
MRR Hits@1 MRR Hits@1 MRR Hits@1
Transductive
NELL995 0.583 0.437 0.578 0.457 0.684 0.555
NELL23k 0.669 0.548 0.756 0.657 0.831 0.762
WDsinger 0.668 0.546 0.720 0.621 0.823 0.738
ConceptNet100k 0.181 0.083 0.650 0.469 0.795 0.658
CoDExSmall 0.900 0.820 0.961 0.935 0.982 0.970
CoDExLarge 0.892 0.824 0.902 0.837 0.973 0.950
YAGO310 0.646 0.403 0.783 0.598 0.971 0.943
AristoV4 0.254 0.201 0.389 0.265 0.597 0.496
DBpedia100k 0.650 0.509 0.717 0.582 0.919 0.861
Hetionet 0.634 0.524 0.809 0.707 0.940 0.890
FB15k-237(10) 0.688 0.550 0.795 0.711 0.918 0.876
FB15k-237(20) 0.695 0.558 0.834 0.758 0.952 0.923
FB15k-237(50) 0.717 0.591 0.876 0.812 0.968 0.946
Table 14: Finetuned inductive relation prediction results (MRR and Hits@1) for Ultra, Trix, and Flock. Bold indicates the best score per row.
Dataset Ultra Trix Flock
MRR Hits@1 MRR Hits@1 MRR Hits@1
Inductive $e,r$
FB-25 0.684 0.563 0.805 0.724 0.909 0.857
FB-50 0.696 0.575 0.780 0.699 0.881 0.820
FB-75 0.754 0.638 0.822 0.699 0.911 0.854
FB-100 0.851 0.769 0.921 0.880 0.965 0.939
WK-25 0.897 0.834 0.905 0.860 0.968 0.954
WK-50 0.865 0.793 0.881 0.840 0.925 0.876
WK-75 0.911 0.875 0.937 0.910 0.965 0.948
WK-100 0.924 0.879 0.916 0.885 0.970 0.946
NL-0 0.632 0.502 0.655 0.518 0.731 0.602
NL-25 0.737 0.622 0.709 0.606 0.757 0.634
NL-50 0.808 0.704 0.774 0.683 0.814 0.721
NL-75 0.795 0.678 0.790 0.671 0.848 0.774
NL-100 0.803 0.678 0.885 0.793 0.937 0.887
MT1 tax 0.990 0.984 0.995 0.990 0.999 0.998
MT1 health 0.929 0.867 0.973 0.949 0.994 0.988
MT2 org 0.981 0.963 0.987 0.978 0.994 0.988
MT2 sci 0.977 0.961 0.990 0.984 0.995 0.992
MT3 art 0.907 0.851 0.887 0.828 0.950 0.916
MT3 infra 0.966 0.947 0.970 0.952 0.996 0.993
MT4 sci 0.954 0.929 0.972 0.952 0.983 0.968
MT4 health 0.951 0.919 0.986 0.979 0.995 0.991
Metafam 0.368 0.112 0.265 0.024 0.997 0.995
FBNELL 0.720 0.576 0.766 0.639 0.879 0.801
Inductive $e$
FB-v1 0.650 0.513 0.705 0.599 0.855 0.766
FB-v2 0.675 0.547 0.713 0.590 0.887 0.812
FB-v3 0.677 0.556 0.742 0.644 0.879 0.810
FB-v4 0.690 0.560 0.766 0.665 0.884 0.807
WN-v1 0.844 0.754 0.776 0.591 0.926 0.879
WN-v2 0.834 0.766 0.765 0.574 0.927 0.869
WN-v3 0.707 0.577 0.756 0.594 0.950 0.911
WN-v4 0.861 0.795 0.804 0.651 0.943 0.898
NL-v1 0.719 0.504 0.590 0.341 0.883 0.766
NL-v2 0.668 0.549 0.811 0.740 0.911 0.870
NL-v3 0.646 0.484 0.757 0.643 0.868 0.795
NL-v4 0.570 0.412 0.822 0.735 0.906 0.849
ILPC Small 0.922 0.876 0.919 0.872 0.953 0.918
ILPC Large 0.875 0.799 0.894 0.829 0.953 0.915
HM 1k 0.626 0.447 0.663 0.414 0.756 0.561
HM 3k 0.592 0.439 0.664 0.418 0.790 0.623
HM 5k 0.605 0.452 0.672 0.428 0.744 0.591
HM Indigo 0.726 0.614 0.835 0.746 0.946 0.903
Table 15: Finetuned transductive relation prediction results (MRR and Hits@1) for Ultra, Trix, and Flock.
Dataset Ultra Trix Flock
MRR Hits@1 MRR Hits@1 MRR Hits@1
Pretraining
FB15k-237 0.795 0.709 0.924 0.870 0.976 0.957
WN18RR 0.914 0.871 0.783 0.634 0.982 0.968
CoDExMedium 0.919 0.870 0.931 0.886 0.974 0.952
Transductive
NELL995 0.630 0.513 0.578 0.457 0.713 0.584
NELL23k 0.688 0.571 0.755 0.658 0.869 0.805
WDsinger 0.730 0.603 0.721 0.627 0.885 0.815
ConceptNet100k 0.612 0.488 0.712 0.551 0.885 0.813
CoDExSmall 0.942 0.900 0.964 0.943 0.981 0.967
CoDExLarge 0.907 0.850 0.908 0.845 0.973 0.950
YAGO310 0.930 0.891 0.826 0.666 0.970 0.942
AristoV4 0.254 0.201 0.498 0.381 0.651 0.547
DBpedia100k 0.650 0.509 0.780 0.665 0.923 0.869
Hetionet 0.737 0.646 0.922 0.862 0.942 0.897
FB15k-237(10) 0.688 0.550 0.795 0.711 0.940 0.905
FB15k-237(20) 0.695 0.558 0.846 0.778 0.958 0.931
FB15k-237(50) 0.728 0.618 0.903 0.858 0.970 0.948
@@ Appendix G Further experimental details
Full results of Section 5.2.
Full tables of zero-shot inductive and transductive entity prediction results for Flock are presented in Table 8 and Table 9, and full tables of finetuned performance are given in Table 10 and Table 11. We further provide the complete zero-shot and finetuned relation prediction results in Table 14, Table 13, and Table 15. Table 19 presents the pretraining graph mix shown in Section 5.3. Finally, detailed hyperparameter settings can be found in Table 20, while the number of epochs used for finetuning is summarized in Table 22.

### v2

@@ Results.
We report the average zero-shot performance for entity prediction and relation prediction over 54 KGs in Tables˜8(a) and 8(b), respectively, as well as trained performance for Petals in Table˜8(c). Across all tasks, Trix with naive noise injection fails to close the gap between Flock. In particular, Trix + noise degrades compared with vanilla Trix without noise injection in relation prediction, while boosting the performance in the entity prediction task. We hypothesize that such a difference lies in the added randomization breaks symmetry among entity embeddings more than among relation embeddings, and entity prediction depends more on having distinguishable entity representations than relation prediction does. Additionally, we attribute this performance gap between Flock and Trix + noise to the source of randomization. Flock introduces stochasticity through random walks, which induces structure-informed perturbations that respect the underlying topology. In contrast, Trix with naive noise injection attempts to break deterministic node-relation equivariance by introducing structure-agonistic noise, which might, in turn, hurt the model’s generalization. Together, these findings suggest that simply adding structure-agonistic noise is insufficient; performance gains only arise when stochasticity is topology-aware and is induced from the graph structure in a principled way.
Appendix G Case study: relation embedding on Metafam
Figure 6: PCA of relation embeddings on Metafam. Ultra (Left) maps several inverse pairs (e.g., fatherOf vs. sonOf) to almost similar embeddings, where Flock (Right) yields clearly separated embeddings, indicating that its probabilistic equivariance allow Flock to distinguish between these semantically different relations, explaining its strong zero-shot performances.
Setup.
To further showcase why expressivity matters for zero-shot generalization, we present a case study on the Metafam dataset (Zhou et al., 2023). Metafam is built from a fixed family-relations ontology: during training, models observe edges with relations motherOf, fatherOf, daughterOf, sonOf, while the test queries only involve motherOf and fatherOf. Up to gender symmetries, this reduces to two effective predictive patterns (parent_of vs. child_of), and the test set focuses on a single one (parent_of).
In the zero-shot setting, KGFMs cannot adapt to this ontology and must rely on their pretrained relation representations. Notice that here, Metafam is challenging: many relations are structurally similar (e.g., fatherOf, sonOf, sisterOf, nieceOf) yet encode opposite predictive patterns.
Result.
Figure˜6 shows that Ultra’s relation embeddings largely collapse these families, placing fatherOf and sonOf in almost identical positions in the PCA plane. This collapse makes it difficult to distinguish who is the parent and who is the child, leading to poor zero-shot performance. Flock, in contrast, can distinguish between these relations even if they are structurally similar, thanks to its random-walk sampling which introduce probabilistic equivariance on nodes and relations. As a results, Flock can produce distinct embeddings to fatherOf and sonOf and achieves much stronger zero-shot performance on Metafam.
Appendix H Analysis of KG sparsity and performance
Figure 7: The zero-shot entity prediction performance of Flock relative to Ultra, plotted against the densities of the 53 KGs. Performance of Flock and log-density of KGs have a Pearson correlation coefficient of -0.53 with p-value 5.0e-5, showing a statistically significant negative correlation.
Setup.
While Appendix˜G explains Flock’s high performance for Metafam, understanding its performances for other general KGs would be beneficial. Thus, we present an additional analysis on the 53 remaining KGs by identifying a structural property that is indicative of the performance of Flock. For the performance measure, we use the relative gain of Flock’s zero-shot entity prediction MRR compared to Ultra. For the structural property, we focus on density of KGs defined by $\frac{|E|}{|V|(|V|-1)}$ which affects the speed of random walks traversing all edges of a KG (intuitively because more edges means more time needed to traverse all of them), and hence is relevant in the context of our theory in Section˜4.2. We make the argument more grounded below.
Result.
Figure˜7 shows that Flock tends to perform well on sparse KGs, while less so on dense KGs, and the tendency is statistically significant. Interestingly, this agrees with our theoretical analysis in Section˜4.2, in which a necessary condition for universality is that the random walk covers all edges of a KG with a high probability (Proposition 4.1). The time taken until covering all edges is called the edge cover time, and it is known to be e.g., $O(|V||E|)$ for uniform walks (Zuckerman, 1991), which is proportional to the density of a graph. This suggests the performance of Flock is associated with the easiness to visit as many edges as possible rapidly, which is more challenging for dense KGs. This analysis is consistent with the observations of recent works on graph learning based on random walks, e.g., Wang & Cho (2024, Section 6) and Kim et al. (2025).
Appendix I Further experimental details
@@
Full results of Section˜5.2.
Full tables of zero-shot entity prediction results are presented in Table˜9, and full tables of finetuned performance are given in Table˜10. We further provide the complete zero-shot and finetuned relation prediction results in Table˜11 and Table˜12. The dataset statistics are in Table˜13, Table˜14 and Table˜15. Table˜16 presents the pretraining graph mix shown in Section˜5.3. Finally, detailed hyperparameter settings can be found in Table˜17, Table˜18 and Table˜19.